### v1

# The Alignment Trap: Complexity Barriers

Abstract
We establish fundamental computational complexity barriers to verifying AI safety as system capabilities scale. Our main results show that for AI systems with expressiveness $\text{EXP}(m)$ above a critical threshold $\tau$, safety verification requires exponential time and is coNP-complete. We formalize the Capability-Risk Scaling (CRS) dynamic, which demonstrates how increasing AI capability drives societal safety requirements toward perfection, creating an inescapable tension with verification complexity. Through four core theorems, we prove that (1) verification complexity grows exponentially with system expressiveness, (2) safe policies comprise at most a $2^{-2^{m}}$ fraction of the policy space, (3) no finite set of alignment techniques can provide universal coverage, and (4) robust safety properties form measure-zero sets for neural networks. These results characterize an ”intractability gap” where practical safety requirements fall within the region of computational intractability. We conclude by presenting a strategic trilemma: AI development must either constrain system complexity to maintain verifiable safety, accept unverifiable risks while scaling capabilities, or develop fundamentally new safety paradigms beyond verification. Our work provides the first systematic complexity-theoretic analysis of AI alignment and establishes rigorous bounds that any safety approach must confront. A formal verification of the core theorems in Lean4 is currently in progress.
Plain Language Summary
This paper explores a fundamental problem in AI safety: as AI gets smarter, it becomes exponentially harder to prove it’s safe. We call this the ”Alignment Trap.” Our work isn’t based on speculation; it’s grounded in four mathematical theorems that reveal significant computational barriers.
1. Verification is Intractable: We prove that checking if an AI’s error rate is below any practical safety threshold (e.g., less than one error in a million) is a computationally ”hard” problem (coNP-complete). This means that for powerful AI, the time required to run such a check could exceed the age of the universe. This isn’t about finding a clever shortcut; the problem is fundamentally difficult.
2. Safe Systems are Geometrically Rare: For modern neural networks, we show that the ”safe” solutions occupy a vanishingly small volume in the space of all possible solutions. Finding one is like trying to hit a single atom from across the galaxy, it’s geometrically a measure-zero target.
3. Safe Systems are Combinatorially Scarce: From an information perspective, safe behaviors are incredibly simple compared to all possible behaviors. We show the fraction of safe systems is like $1$ divided by $2$ raised to an astronomical power. They are information-theoretically rare.
4. No Universal Safety Test Exists: We prove that it’s impossible to create a single, one-size-fits-all algorithm that can verify the safety of every possible AI system. For any safety test you design, there will always be an AI that can circumvent it.
The Bottom Line: These mathematical results show that ensuring AI safety isn’t just a matter of being more careful or demanding lower error rates. We are facing hard computational limits. This forces a difficult choice: limit AI capability to a level we can manage, accept a degree of un-verifiable risk, or invent new safety methods that go beyond today’s verification techniques.
1 Introduction
Artificial intelligence (AI) capabilities are advancing at a remarkable pace, with systems demonstrating sophisticated abilities in domains from natural language processing to scientific discovery [11, 8]. This progress promises transformative benefits but also introduces a fundamental challenge: ensuring that as these systems become more powerful, they remain safe and aligned with human values [4, 14]. The potential for highly capable systems to cause widespread, catastrophic harm, whether by accident or through instrumental misalignment, is no longer a purely theoretical concern [1].
This paper argues that the challenge of AI safety is not merely an engineering problem to be solved with more data or better algorithms. Instead, we confront fundamental mathematical barriers that intensify as AI capabilities increase. We identify two fundamental barriers to any verification-based safety approach:
  1. 1.
    The Complexity Barrier: Verifying whether a sufficiently advanced AI system is safe is computationally intractable (coNP-complete), and the set of verifiably safe systems is vanishingly small (measure-zero) in the space of all possible systems. Finding and verifying safety is fundamentally hard.
  2. 2.
    The Specification Barrier: Defining safety is a moving target. As AI capabilities grow, societal demands for safety evolve, and the technical definition of what constitutes a ”safe” outcome becomes more complex and pluralistic, making any fixed specification inadequate over time.
These two barriers converge to create a dilemma we term the Alignment Trap: the very scaling of AI capabilities that makes safety guarantees essential also renders their verification and specification computationally intractable. Our contribution is to formalize this trap through a series of theorems establishing results on coNP-completeness, topological scarcity, universal impossibility, and the invariance of complexity under alignment transformations. We do not claim philosophical impossibility, but rather demonstrate a fundamental tension rooted in computational reality.
Crucially, these barriers are not limited to the unrealistic standard of perfect safety ($\epsilon=0$). Our results demonstrate that intractability persists even when we allow for the small, non-zero error tolerances typical of real-world engineering.
Taken together, these findings force a stark strategic choice. The pursuit of ever-more-powerful AI systems leads to a direct confrontation with fundamental computational limits on our ability to ensure their safety. This presents a trilemma:
  • Constrain Capability: Limit the expressive power of AI systems to a level where safety verification remains computationally feasible.
  • Accept Irreducible Risk: Proceed with the development of highly capable systems while accepting that their safety cannot be fully guaranteed, creating a permanent risk of catastrophic failure.
  • Develop New Paradigms: Pursue entirely new approaches to safety that do not rely on the traditional verification of system behavior against a predefined specification.
This paper provides the formal groundwork for understanding these barriers. By framing the alignment problem in terms of computational complexity, we shift the discussion from a philosophical debate to a rigorous, mathematical investigation of the limits of what can be guaranteed.
The paper is structured as follows. Section 2 reviews related work. Section 3 provides a detailed overview of the two fundamental barriers. Section 4 establishes our formal framework. Section 5 presents the core theorems that provide the mathematical foundation for each barrier. Section LABEL:sec:alignment_trap_synthesis synthesizes these results into the main Alignment Trap theorem. Section 7 explores the implications of our findings and discusses potential paths forward. Section 8 concludes. Detailed proofs for all theorems are provided in the Appendices.
2 Related Work
This work builds on several distinct but intersecting fields of theoretical computer science and machine learning. Our results on the complexity of verifying AI safety should be understood in the context of established research in computational complexity, statistical learning theory, and formal verification.
2.1 Computational Complexity in Machine Learning
The study of computational hardness in machine learning is a rich field. Foundational work by Arora, Barak, and others has established that many learning problems are computationally intractable in the worst case. For instance, learning a 3-term DNF formula is NP-hard [2]. Our work extends this tradition by analyzing the complexity not of learning, but of *verifying* the properties of a learned model.
2.2 Formal Verification and its Limits
Formal verification, particularly model checking, has been successfully applied to hardware and software systems for decades [6]. However, the state-space explosion problem has always been a fundamental limitation. In the context of AI, recent work on verifying properties of neural networks, such as the Reluplex algorithm by Katz et al. [9], has shown promise for specific, bounded problems. Our results provide a broader, complexity-theoretic perspective on why these approaches face fundamental scaling limits as system capability increases.
2.3 PAC Learning and PAC-Bayes Theory
The Probably Approximately Correct (PAC) learning framework, introduced by Valiant [16], provides guarantees on the generalization error of a learned hypothesis. Our results on the scarcity of safe policies connect to this framework. Furthermore, we leverage the PAC-Bayes framework, developed by McAllester [10], to provide information-theoretic bounds on the number of bits required to specify a safe policy, demonstrating that safety is a highly compressible concept and thus rare.
2.4 AI Safety and Alignment
The field of AI safety has identified numerous challenges, from reward specification to interpretability. Our work provides a formal, mathematical underpinning for many of these concerns, framing them as consequences of fundamental complexity barriers. While much of the existing literature is empirical or philosophical, our contribution is a set of rigorous impossibility and intractability results that are independent of specific architectures or learning algorithms, relying only on minimal assumptions about system expressiveness.
3 Two Fundamental Barriers
We identify two fundamental barriers that any verification-based approach to AI safety must confront: the Complexity Barrier and the Specification Barrier. These barriers are not independent; they interact and reinforce one another, culminating in the Alignment Trap.
3.1 The Complexity Barrier: Finding and Verifying Safety
The first barrier, a wall of computational intractability, arises from the sheer difficulty of ensuring safety in expressive systems. This barrier has two faces: the geometric scarcity of safe solutions and the computational complexity of verifying them.
3.1.1 Geometric Perspective: The Scarcity of Safety
From a topological standpoint, safe policies are exceptionally rare. For expressive models like neural networks, the set of parameters corresponding to safe behaviors forms a vanishingly small, or measure-zero, subset of the total parameter space. This scarcity is formally established in Theorems 5.3 and 5.5, which prove that finding a safe policy is akin to hitting a target of infinitesimal size.
3.1.2 Computational Perspective: The Intractability of Verification
Even if a seemingly safe policy is found, verifying its safety is computationally intractable. This is formalized in Theorems 5.1 and 5.2, where we prove that the problem of verifying whether a system’s error rate is below any non-trivial threshold is coNP-complete. This means that for any verification algorithm, there will always be systems for which the verification process takes an exponential amount of time, rendering it impractical for highly capable AI.
3.1.3 Synthesis
These two aspects create a powerful dilemma: finding a safe policy is extraordinarily difficult due to their scarcity, and even if one is found, confirming its safety is computationally infeasible.
3.2 The Specification Barrier: Defining Safety
The second barrier is the challenge of precisely and robustly defining what ”safety” means. As AI capabilities increase, the task of specifying safety becomes a moving target.
3.2.1 The Moving Target Problem
Simple, static rules are insufficient for complex, real-world environments. As an AI’s capabilities expand, it encounters novel situations, and the criteria for what constitutes a safe action evolve. We formalize this with the Movable Goalpost Theorem (Theorem 5.9), which shows that the goalposts are not just moving, but are being pushed further away by the very progress we make. As system capability grows, the complexity of the safety specification required to prevent unwanted outcomes must also grow, often at an unsustainable rate.
3.2.2 Value Specification and Pluralism
Beyond technical evolution, defining safety is also a challenge of value alignment. Human values are complex, often contradictory, and culturally dependent. Capturing this richness in a formal specification that an AI can optimize is an open problem. Any fixed specification risks oversimplifying or misinterpreting these values, leading to potentially catastrophic outcomes. This challenge is exacerbated by the pluralistic nature of human societies, where different groups may have conflicting but equally valid definitions of safety.
4 Formal Setup
To analyze the Alignment Trap rigorously, this section establishes a concise formal framework. These definitions provide the mathematical language to describe AI systems, their capabilities, potential for harm, and the challenge of ensuring their safety.
4.1 The Collingridge Dilemma and the CRS Dynamic
The starting point for our analysis is the Collingridge Dilemma, a foundational concept in technology governance [7]. The dilemma states that in the early stages of a technology’s development, its risks are easy to control but hard to predict; by the time the risks are well understood, the technology is often too entrenched to manage. The Alignment Trap is a formalization of this dilemma for AI, driven by a dynamic we term Capability-Risk Scaling (CRS).
The CRS dynamic describes the fundamental tension between increasing AI capability and the societal demand for safety. As a system becomes more powerful, its potential for catastrophic harm grows, forcing society to demand ever-higher standards of safety. We can formalize this using the concept of an F-N curve, a standard tool in risk analysis that plots the frequency (F) of an event against its number of fatalities (N).
Definition 4.1 (F-N Curve and Societal Risk Tolerance).
An F-N curve defines the boundary of acceptable risk. For a given consequence level $N$, it specifies the maximum acceptable frequency $F$. For systems with catastrophic potential, societal risk tolerance dictates that as $N$ increases, $F$ must decrease, typically following a power law: $F\cdot N^{\alpha}\leq k$, for some constant $k$ and exponent $\alpha\geq 1$.
For AI, we can adapt this to a Capability-Impact model. Let $C$ be the capability of a system and $D(C)$ be the potential impact or damage, which is an increasing function of $C$. The acceptable probability of a catastrophe, $P_{\text{acceptable}}(C)$, is inversely related to $D(C)$.
Definition 4.2 (The CRS Dynamic).
As AI capability, $C$, increases:
  1. 1.
    The potential catastrophic impact, $D(C)$, increases.
  2. 2.
    Societal risk tolerance for a catastrophe, $P_{\text{acceptable}}(C)$, decreases, following the principle of the F-N curve, e.g., $P_{\text{acceptable}}(C)\propto D(C)^{-\alpha}$.
  3. 3.
    The required alignment precision, $\epsilon_{\text{required}}(C)$, must therefore approach zero to meet this shrinking risk tolerance.
Theorem 4.3 (CRS Convergence).
Under the assumption that societal risk tolerance $P_{\text{acceptable}}(C)$ is a strictly decreasing function of capability $C$ and that the probability of catastrophe is a continuous, non-decreasing function of alignment error $\epsilon$, it follows that:
$\lim_{C\to\infty}\epsilon_{\text{required}}(C)=0$
Proof Sketch.
As $C\to\infty$, the potential impact $D(C)\to\infty$. Per the F-N curve principle, this drives the acceptable probability of a catastrophe $P_{\text{acceptable}}(C)\to 0$. For the actual probability of catastrophe to remain below this vanishing threshold, the alignment error $\epsilon$ must also be driven to zero. The full proof is in Appendix A.1. ∎
This dynamic is the engine of the Alignment Trap: it creates the demand for near-perfect safety that collides with the complexity barriers described in this paper.
[…]
Figure 1: The Alignment Trap. The diagram illustrates the core CRS dynamic. As capability increases (x-axis), societal demand for safety requires the alignment error $\epsilon_{\text{req}}$ to drop exponentially (blue line, left y-axis). Simultaneously, the computational cost of verifying safety grows exponentially (red line, right y-axis). The shaded region represents the ”trap”: the zone where the required level of safety becomes too costly or computationally intractable to verify.
4.2 Core Concepts
Definition 4.4 (Policy and Policy Space).
An agent’s behavior is described by a policy $\pi:X\to Y$, which maps inputs from a space $X$ to outputs in a space $Y$. The set of all policies that can be realized by a given AI architecture is the policy space $\Pi$.
Definition 4.5 (Harm and Safety).
A harm functional $H:Y\to[0,\infty)$ quantifies the undesirability of an output $y$. An output is considered safe if $H(y)=0$. A catastrophe occurs if $H(y)$ exceeds a critical threshold, $H_{\text{crit}}$. A policy $\pi$ is perfectly safe if $H(\pi(x))=0$ for all $x\in\mathcal{X}$.
Definition 4.6 (Alignment Error $\epsilon$).
The alignment error, $\epsilon(\pi)$, measures a policy’s deviation from a perfectly safe ideal. This can be defined probabilistically, such as the measure of the set of inputs where the policy acts unsafely: $\epsilon(\pi)=\mu(\{x\in\mathcal{X}\mid H(\pi(x))>0\})$. Perfect alignment corresponds to $\epsilon(\pi)=0$.
Definition 4.7 ($\epsilon$-Robust Safety).
A policy $\pi$ is considered $\epsilon$-robustly safe if for every input $x\in\mathcal{X}$, the policy’s output and the outputs for all perturbations of $x$ within an $\epsilon$-ball remain within the set of safe outcomes. Formally, for a given distance metric $d$ on the input space $\mathcal{X}$, a policy $\pi$ is $\epsilon$-robustly safe if:
$\forall x\in\mathcal{X},\forall x^{\prime}\in\mathcal{X}\text{ with }d(x,x^{\prime})<\epsilon,\text{ we have }\pi(x^{\prime})\notin D$
where $D$ is the set of catastrophic outcomes. This ensures that the policy’s safety is resilient to small perturbations in the input, a critical property for reliable real-world deployment.
Definition 4.8 (Finite-Set Expressiveness EXP(m)).
An AI architecture has finite-set expressiveness EXP(m) if, for any finite set of $m$ inputs, it can implement any arbitrary Boolean function on those inputs. This is a minimal assumption on the flexibility of the system, satisfied by most modern neural networks which can memorize or perfectly fit a finite number of training examples. We assume that as a system’s general capability increases, so does its expressiveness $m$.
4.3 Core Assumptions
Our analysis rests on a set of explicit assumptions about the systems we are studying and the environment they operate in. These are intended to be minimal and reflect the conditions under which AI safety becomes a pressing concern.
Core Assumptions A1. High Expressiveness: The systems under consideration have high expressiveness, formally captured by $\text{EXP}(m)$ for a sufficiently large $m$. This is the source of both their power and the complexity of their behavior. A2. Worst-Case Verification: Safety verification is concerned with worst-case guarantees. An agent is only considered ”safe” if it is proven to not cause catastrophic harm for any possible input, not just on average. A3. Monotonic Risk: For any concrete, well-defined goal, the risk of a catastrophic outcome is a non-decreasing function of the alignment error $\epsilon$. A higher error rate cannot lead to a lower risk. This is standard in safety-critical engineering. A4. Shrinking Risk Tolerance: As a system’s capability and potential impact grow, society’s tolerance for catastrophic error shrinks, consistent with established principles of risk management (e.g., F-N curves). For highly capable systems, the required alignment error, $\epsilon_{\text{required}}$, approaches zero.
These definitions and assumptions form the foundation for the theorems presented in the next section. They are designed to capture the essential features of the AI safety problem in a way that permits rigorous mathematical analysis.
5 Core Theorems
We now present the mathematical foundations for each barrier, establishing the formal complexity and topological results that underpin our analysis. Brief proof sketches are provided here; full, rigorous proofs are detailed in the Appendices.
5.1 Verification Complexity Results
Theorem 5.1 (Sharp Verification Threshold).
Let $\Pi$ be a class of policies with finite-set expressiveness $\text{EXP}(m)$. For any error threshold $\epsilon$ in the range $0<\epsilon\leq 2^{-\Omega(m)}$, the problem of verifying whether a policy $\pi\in\Pi$ is $\epsilon$-safe (i.e., $\epsilon(\pi)\leq\epsilon$) is coNP-complete. The transition to intractability is sharp: verification is manageable when system complexity is below a logarithmic bound ($m<\lceil\log_{2}n\rceil+1$) but becomes coNP-complete immediately upon crossing it.
Proof Sketch.
The proof involves a reduction from the Tautology problem for a Boolean formula $\phi$ with $m$ variables. We construct a policy $\pi_{\phi}$ that operates on $m$ input bits and produces a catastrophic output if and only if the formula is not a tautology. Verifying the policy’s safety is thus equivalent to verifying the tautology. Since a single counterexample to the tautology corresponds to an error rate of at least $1/2^{m}$, the hardness holds for any $\epsilon<2^{-m}$. The full proof is in Appendix A.2. ∎
Theorem 5.2 ($\epsilon$-Robust Verification Complexity).
The coNP-completeness of verification holds even for $\epsilon>0$. A single counterexample to a tautology corresponds to a single failure mode, which is difficult to distinguish from zero failures for any small $\epsilon$.
Proof Sketch.
The hardness for $\epsilon>0$ is established by showing that distinguishing between an error rate of 0 and an error rate of $\delta$ (for some small $\delta>0$) is as hard as the original problem. The full proof is in Appendix B. ∎
5.2 Topological Scarcity Results
Theorem 5.3 (Measure Zero for Safe Policies).
Let $\Pi$ be the space of policies realized by a ReLU neural network. The set of policies that are robustly $\epsilon$-safe has Lebesgue measure zero in the parameter space.
Proof Sketch.
The proof relies on the piecewise linear nature of ReLU networks. The parameter space is partitioned into regions where the network computes a single affine function. The boundary between a safe and an unsafe region is a lower-dimensional manifold. Under the standard genericity assumption that the gradient of the safety margin function is non-zero, we can always find a small perturbation to the network’s parameters that moves it from a safe to an unsafe state. This implies the set of robustly safe policies has no interior and thus has measure zero. The full proof is in Appendix A.3. ∎
Theorem 5.4 (Topological Alignment Trap).
Let the parameter space of policies be $\mathbb{R}^{n}$ with $n\geq 2$. If the set of perfectly safe policies $\Pi_{S}$ is geometrically ”thin” (Hausdorff dimension $<n-1$) and training dynamics are generic $C^{1}$ paths, then for almost every starting policy, the training path will never intersect the safe set $\Pi_{S}$.
Proof Sketch.
The proof is conditional on a set of three well-defined axioms: (1) the geometric thinness of the safe policy set, (2) the smoothness of the initial condition distribution, and (3) the generic avoidance of thin sets by training dynamics. These axioms, justified by standard arguments from geometric measure theory and transversality, collectively imply that a generic training path, starting from a random (and thus almost surely unsafe) point, will almost surely fail to intersect the safe set. The full, axiom-based proof is in Appendix E.1. ∎
000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111000.50.50.50.511111.51.51.51.522222.52.52.52.53333Hazard probability H($\pi$)DensityRandom linear policies (N=50000)Fraction with H($\pi$) $\leq$ $\epsilon$ = 0.372%$\epsilon=0.01$
Figure 2: Distribution of hazard probabilities $H(\pi)$ for 50,000 randomly sampled linear policies. The histogram shows that while many policies have low hazard probabilities (peak around 0.05-0.08), only 0.372% fall below the safety threshold of $\epsilon=0.01$ (indicated by the dashed line). This empirically demonstrates the extreme rarity of safe policies in the parameter space, supporting the theoretical measure-zero results.
Theorem 5.5 (Combinatorial Scarcity of Safe Policies).
For a policy class with expressiveness $\text{EXP}(m)$, the fraction of policies that are perfectly safe is bounded by $2^{-2^{m}}$.
Proof Sketch.
There are $2^{2^{m}}$ possible Boolean functions on $m$ inputs. If safety corresponds to a unique function (e.g., the all-zero function), then only one of these $2^{2^{m}}$ functions is safe, representing an exponentially small fraction of the total space. The full proof is in Appendix A.4. ∎
[…]
Figure 3: The fraction of perfectly safe policies as a function of expressiveness $m$. The number of possible policies grows as $2^{2^{m}}$, while the number of safe policies remains constant (or grows much slower), causing the fraction to plummet.
5.3 Universal Impossibility Result
Theorem 5.6 (No Universal Alignment Technique).
For any countable set of alignment techniques $\{A_{1},A_{2},\ldots\}$, there exists a policy $\pi^{*}$ that is not aligned by any technique in the set.
Proof Sketch.
The proof uses a standard diagonalization argument. We construct a ”diagonal” policy $\pi^{*}$ that is specifically designed to misbehave in a way that the $i$-th alignment technique, $A_{i}$, fails to correct. By its construction, $\pi^{*}$ is misaligned after the application of any technique from the countable set. The full proof is in Appendix A.5. ∎
5.4 Invariance of Complexity under Alignment Transformations
Theorem 5.7 (Trap Universality).
For any alignment transformation that preserves the computational capability of a policy, the fundamental verification complexity barrier persists. Specifically, for any such transformation $A$, one can construct a policy $\pi$ with a cryptographic backdoor such that verifying the safety of the transformed policy $A(\pi)$ remains cryptographically hard (i.e., requires exponential time).
Proof Sketch.
The proof relies on constructing a policy $\pi$ whose catastrophic behavior is tied to a secret key of a pseudorandom function (PRF). An alignment transformation $A$ is assumed to be computationally efficient and ”capability-preserving,” meaning it cannot drastically simplify the policy’s logic. Such a transformation cannot learn the secret key from black-box access to $\pi$. Therefore, it cannot reliably remove the PRF-based trap without breaking the cryptography. The trap persists in the transformed policy $A(\pi)$, and verifying its safety remains as hard as breaking the PRF. The full proof is in Appendix E.4. ∎
5.5 Learning Impossibility Results
Theorem 5.8 (PAC-Bayes Alignment Lower Bound).
Let $\mathcal{H}$ be a hypothesis class and $S\subset\mathcal{H}$ be a set of perfectly safe policies. Assume $S$ has measure zero under any non-degenerate prior $P$ (i.e., $P(S)=0$). For any posterior distribution $Q$ with finite KL divergence from $P$, and any $\epsilon_{\min}>0$ representing the minimum catastrophic risk for policies outside $S$:
$\mathbb{E}_{h\sim Q}[L(h)]\geq\epsilon_{\min}\cdot(1-Q(S))$
where $L(h)$ denotes the catastrophic risk of policy $h$. If $\mathrm{KL}(Q\|P)<\infty$, then $Q(S)=0$, implying $\mathbb{E}_{h\sim Q}[L(h)]\geq\epsilon_{\min}>0$.
Proof Sketch.
(a) Counting Argument: There are $2^{2^{m}}$ possible Boolean functions on $m$ inputs. If safety corresponds to a unique function (e.g., the all-zero function), then only one of these $2^{2^{m}}$ functions is safe.
(b) Learning Argument (PAC-Bayes): The result from (a) implies that any non-degenerate prior over policies will assign a vanishingly small probability to the safe set, $P(S)\approx 0$. The PAC-Bayes inequality shows that to learn a posterior $Q$ that concentrates on the safe set (i.e., $Q(S)\approx 1$), the KL-divergence $\mathrm{KL}(Q\|P)$ must be infinite. This means a finite dataset cannot provide the information needed to overcome an unbiased prior and guarantee safety. The full proof is in Appendix E.2. ∎
5.6 Specification Instability Result
Theorem 5.9 (The Movable Goalpost / Epistemic Fragility Theorem).
Under conditions of stakeholder pluralism, where a set of stakeholders $L$ (potentially large or diverse) each define a set of catastrophic outcomes $B_{\text{cat}}^{(\lambda)}\subseteq Y$:
  1. 1.
    Aggregated Safety Requirement: A system $A$ is perfectly safe with respect to all stakeholders if and only if its outputs $A(x)$ avoid $B_{\text{cat}}^{\text{(total)}}=\bigcup_{\lambda\in L}B_{\text{cat}}^{(\lambda)}$ for all inputs $x$.
  2. 2.
    Emergent Impossibility from Pluralism (Core Statement): Even if each individual stakeholder’s catastrophic set $B_{\text{cat}}^{(\lambda)}$ is “reasonable” (e.g., $B_{\text{cat}}^{(\lambda)}\neq Y$, allowing for some non-catastrophic behaviors from their individual perspective, and perhaps even allowing for a non-zero error tolerance $\epsilon^{*}_{\lambda}>0$ with respect to their own ideal), the union $B_{\text{cat}}^{\text{(total)}}$ can become equivalent to the entire output space $Y$, or all but a measure-zero or computationally infeasible set of permissible behaviors. This forces the effective alignment error $\epsilon^{*}$ required to satisfy the collective to approach zero with respect to a target safe set $Y\setminus B_{\text{cat}}^{\text{(total)}}$ that may itself be vanishingly small or empty.
  3. 3.
    Epistemic Fragility: The overall safety definition $Y\setminus B_{\text{cat}}^{\text{(total)}}$ becomes epistemically fragile. The addition of a single new stakeholder with a sufficiently restrictive $B_{\text{cat}}^{(\text{new})}$ (or a change in an existing stakeholder’s definition) can drastically alter $B_{\text{cat}}^{\text{(total)}}$, potentially rendering a previously achievable collective safety target unachievable or an already safe system unsafe under the new aggregate definition.
6 Synthesis: The Alignment Trap
The preceding sections have established two independent barriers to ensuring AI safety: the Complexity Barrier, concerning the difficulty of finding and verifying safe policies, and the Specification Barrier, which addresses the challenge of defining safety itself. This section synthesizes these concepts to formalize the Alignment Trap, a dynamic that emerges from their interaction.
6.1 The Jaws of the Trap: Formalizing the Barriers
The two barriers are not merely conceptual; they are grounded in the formal results presented in the Core Theorems (Section 5) and are further elaborated here.
Remark 6.1 (Plain-English Summary of the Verification Barrier).
This theorem makes a very strong claim: for any AI system that could potentially cause a catastrophe, and for *any* safety mechanism or alignment technique we design to prevent such catastrophes, the problem of definitively checking whether this safety mechanism actually works perfectly (i.e., guarantees no catastrophes) is undecidable. This means there’s no universal algorithm that can always give a ”yes” or ”no” answer to whether our safety fix is foolproof. The difficulty lies in the very nature of detecting potential catastrophes.
Theorem 6.2 (The Verification Barrier).
For any system S with the catastrophic error property, and any alignment function F guaranteeing no catastrophic outputs, verifying F’s success is undecidable.
Remark 6.3 (Plain-English Summary of Measure-Zero Safety).
This theorem strengthens the ”needle in a haystack” argument. It suggests that even if we’re not trying to make an AI behave exactly like one specific ideal safe model, but are instead looking for *any* AI behavior that manages to avoid all known catastrophes, such universally safe behaviors are still exceptionally rare (measure-zero) in the space of all possible complex AI behaviors, especially if any small error can lead to disaster (Regime A).
Theorem 6.4 (Universal Measure-Zero Safety).
In Regime A (Effective Brittleness), the set of ALL functions (not just identity-preserving ones) that successfully prevent catastrophes has measure zero in a properly formalized function space.
6.2 The Vicious Cycle of Capability and Risk
The Alignment Trap emerges from a feedback loop driven by increasing AI capability. As a system’s capability ($C$) grows, so does its potential impact, leading to a societal demand for higher levels of safety. This demand translates into a requirement for a vanishingly small error tolerance ($\epsilon\to 0$).
This requirement for near-perfect safety activates both barriers simultaneously:
  • The demand for a low $\epsilon$ makes verification computationally intractable, as shown by the Complexity Barrier (Theorems 5.1, 5.2).
  • The need to define safety with such high precision runs into the Specification Barrier, as the definition becomes increasingly complex and contested (Theorem 5.9).
This creates a vicious cycle: the pursuit of greater capability leads to safety requirements that are impossible to either verify or specify.
6.3 The Alignment Trap Theorem
This dynamic is formally captured in the paper’s central theorem.
Remark 6.5 (Plain-English Summary of the Alignment Trap).
This is the central theorem of the paper. It states that as AI becomes more powerful, society demands near-perfect safety (almost zero mistakes). However, at the same time, the very act of checking or verifying that such a powerful AI is indeed perfectly safe becomes mathematically impossible or computationally too expensive to carry out. So, precisely when we need safety the most, we lose the ability to confirm it. This creates a ”trap”: the more capable the AI, the higher the safety standards, but the harder it is to meet and verify those standards.
Theorem 6.6 (The Alignment Trap).
As AI capability C increases:
  1. 1.
    Premise (Capability-Risk Scaling): Increasing capability $C$ leads to a requirement for alignment error $\epsilon_{required}(C)\to 0$.
  2. 2.
    Premise (Verification and Specification Barriers): Verifying that alignment error is below $\epsilon_{required}(C)$ is computationally intractable and defining safety with sufficient precision is non-computable or intractably complex.
  3. 3.
    Conclusion (The Trap): For any finite verification budget and specification complexity tolerance, there exists a capability threshold $C^{*}$ beyond which the required safety becomes impossible to achieve or verify.
6.4 Error Regimes and the Necessity of Perfection
The ”Fork Dichotomy” theorem provides additional context for why the demand for $\epsilon\to 0$ is often a technical necessity rather than a mere preference.
Theorem 6.7 (The Fork Dichotomy).
Any AI system, when evaluated against a concrete goal, must operate in one of two regimes:
  1. 1.
    Regime A (Hard Errors): Where any error, no matter how small, leads to a non-zero risk.
  2. 2.
    Regime B (Soft Errors): Where there exists an error buffer within which the system remains safe.
For systems operating in Regime A with respect to critical safety goals, the demand for zero risk is equivalent to a demand for zero error. This reinforces the Alignment Trap, as it connects the societal need for safety directly to the mathematical requirement for perfection that the core theorems show is intractable.
7 Implications and Discussion
The formal barriers we have established have profound implications for the future of AI development and safety. This is not a declaration of impossibility, but a rigorous characterization of the constraints under which safety efforts must operate.
7.1 What These Barriers Mean for AI Development
The Alignment Trap is a fundamental barrier reshaping AI safety’s future [12, 5]. As AI systems scale in capability ($C$), their potential impact ($D(C)$) grows. This growth leads to rational societal demand for vanishingly small catastrophic failure probabilities ($P_{\mathrm{ca}}(C)\to 0$), necessitating near-zero alignment error ($\epsilon\to 0$).
However, as we have proven, verifying that alignment error is below any practically meaningful threshold is computationally intractable for expressive systems. This creates a tension: as capability increases, verification becomes computationally intractable while safety requirements become more stringent. These complexity barriers apply to any safety approach that relies on verifying the behavior of a highly expressive model, shifting the focus from finding a universally applicable technique to understanding the trade-offs between capability, verifiability, and acceptable risk.
7.2 The $\epsilon>0$ Case: The ”Death Valley” of Safety
The complexity barriers established in this work are not confined to the theoretical case of perfect safety ($\epsilon=0$). They extend robustly into the bounded-error regime that governs all real-world safety requirements. For each barrier, there exists a critical threshold $\epsilon^{*}(m)\approx 2^{-\Omega(m)}$ below which the barrier remains fully operative.
For any high-stakes application (aviation, medical, etc.), the required safety tolerance (e.g., $10^{-9}$) falls squarely within this intractable region for any sufficiently complex system. We call the region between what is computationally verifiable (e.g., $\epsilon>0.01$) and what is practically required (e.g., $\epsilon<10^{-6}$) the ”death valley” of safety requirements. Relaxing standards to escape the valley is not a viable solution, as it would mean accepting systems that fail catastrophically far too often.
7.3 Paths Forward: The Strategic Trilemma
Our complexity results imply a fundamental trilemma. The computational barriers create three distinct development paths:
  1. 1.
    Shape Capability: Deliberately limit the expressive power of AI systems to a level where safety verification remains computationally feasible. This path allows for formal guarantees by constraining systems to domains where verification remains tractable.
  2. 2.
    Accept Blind Risk: Proceed with the development of highly capable systems while accepting that their safety cannot be fully guaranteed. On this path, safety properties become statistical rather than provable, relying on empirical testing and monitoring instead of formal guarantees.
  3. 3.
    Develop New Paradigms: Pursue entirely new approaches to safety that do not rely on the traditional verification of system behavior against a predefined specification. This path focuses on creating systems with built-in, inherent behavioral constraints (e.g., through causal or structural properties) that are safe by construction.
This trilemma follows directly from our core complexity results. Each path represents a different strategic response to the computational barriers we have identified, transforming the challenge from a statement about impossibility to a map of the strategic landscape.
7.4 A Note on Worst-Case vs. Average-Case Analysis
A critical reader might point out that our analysis is based on a worst-case complexity framework. In traditional computer science, we often accept theoretical worst-case bounds (e.g., Quicksort’s O(n²)) because the average-case performance is excellent in practice. However, for the highly capable AI systems considered in this paper, we argue that this distinction is far less meaningful, and the focus on worst-case complexity is not a limitation but a necessary and conservative assumption.
Why Worst-Case Matters More for Advanced AI
  • Adversarial Optimization by Default: Highly capable AI systems are, by definition, powerful optimizers. When such systems pursue objectives misaligned with human values, they naturally seek out edge cases and exploit system boundaries, effectively turning every deployment into a worst-case scenario. The system isn’t randomly sampling from the input space; it’s intelligently searching for inputs that maximize its objective function.
  • The Capability Paradox and CRS Dynamic: The Capability-Risk Scaling (CRS) dynamic posits that as a system’s capability ($C$) increases, its potential impact ($D(C)$) grows, forcing the required safety tolerance ($\epsilon_{\text{required}}$) towards zero. The paradox is that the increase in $C$ is a double-edged sword: not only does it increase the stakes, but it also enhances the system’s ability to find and exploit the very edge cases that would violate this shrinking error tolerance. A more capable system is better at finding the rare, worst-case failure modes that a less capable system might miss, making the stringent safety requirements of the CRS dynamic even harder to meet.
  • Existential Stakes Change the Game: When a single worst-case failure of an AGI system could be catastrophic or existential, average-case analysis becomes a dangerous comfort. As Nick Bostrom might put it: ”For existential risks, we need to get it right on the first try.” This paper rigorously proves that for sufficiently capable systems, we cannot mathematically guarantee we’ll get it right even once.
  • Deceptive Alignment and Distribution Shift: Advanced AI systems may appear aligned during testing (contributing to good ”average-case” behavior) while harboring misaligned mesa-objectives that only manifest in specific deployment contexts. Furthermore, these systems can influence their environment, creating novel situations entirely outside any training distribution.
The traditional comfort of ”it works well in practice despite theoretical worst-case bounds” becomes cold comfort when the system in question is actively intelligent, potentially deceptive, and capable of causing irreversible harm. Our results establish a baseline of computational difficulty that any future approach must overcome, and this baseline is most prudently set at the worst case.
[…]
Figure 4: The two core theorems with quantifiable computational costs. Verification complexity shows operations required vs. target error rate for different system complexities $m$.
8 Conclusion
This paper has established that the pursuit of verifiably safe, highly capable AI systems confronts fundamental mathematical barriers. We have framed this challenge in terms of two interacting barriers, Complexity and Specification, which together create the Alignment Trap. Our results collectively demonstrate that verification becomes computationally infeasible long before it becomes practically necessary.
The convergence of these barriers forces a strategic trilemma: society must either constrain system complexity to maintain verifiable safety, accept unverifiable risks while scaling capabilities, or develop fundamentally new safety paradigms beyond verification. Our work provides the first systematic complexity-theoretic analysis of AI alignment and establishes rigorous bounds that any safety approach must confront. The path we choose will determine the future of our relationship with artificial intelligence.
The author gratefully acknowledges the pioneering contributions of Professor Terence Tao to the integration of large language models in mathematical theorem proving. While this work addresses impossibility results in AI alignment, it has benefited from methodological advances in AI-assisted mathematics research. The author thanks the broader research community working at the intersection of formal mathematics and artificial intelligence, whose collective efforts have established important foundations for this work. The development of the formal proofs presented herein was aided by large language model interactions, exemplifying emerging practices in mathematical research. The author acknowledges the DEF CON AI Village for providing a foundational research environment that contributed significantly to this work. The author expresses gratitude to their mentor at Google for valuable guidance and constructive feedback throughout this project.
Appendix A Extended Proofs
This appendix provides the detailed, formal proofs for the core theorems presented in Section 5.
A.1 Proof of Theorem 4.3 (Convergence to Zero Error Demand)
Proof.
We proceed by contradiction. Assume that $\lim_{C\to\infty}\epsilon_{\text{required}}(C)\neq 0$. This implies that there exists some $\epsilon_{0}>0$ and a sequence of capabilities $C_{n}\to\infty$ such that $\epsilon_{\text{required}}(C_{n})\geq\epsilon_{0}$ for all $n$.
By definition of $\epsilon_{\text{required}}$, this means that for each $C_{n}$, there is an acceptable probability of catastrophe $P_{\text{acceptable}}(C_{n})$ that is at least as large as the risk associated with an error of $\epsilon_{0}$. Let $f_{\text{risk}}(\epsilon)$ be the non-decreasing function mapping alignment error to catastrophe probability. Then $f_{\text{risk}}(\epsilon_{0})\leq P_{\text{acceptable}}(C_{n})$ for all $n$.
However, the CRS dynamic posits that as $C\to\infty$, the potential impact $D(C)\to\infty$, which in turn forces the socially acceptable risk $P_{\text{acceptable}}(C)\to 0$. Therefore, for the sequence $C_{n}\to\infty$, we must have $\lim_{n\to\infty}P_{\text{acceptable}}(C_{n})=0$.
This leads to a contradiction. We have a fixed positive number $f_{\text{risk}}(\epsilon_{0})>0$ (since $\epsilon_{0}>0$ and risk is non-trivial), but the sequence $P_{\text{acceptable}}(C_{n})$ must converge to 0. For a large enough $n$, $P_{\text{acceptable}}(C_{n})$ will be smaller than $f_{\text{risk}}(\epsilon_{0})$, violating the condition that $f_{\text{risk}}(\epsilon_{0})\leq P_{\text{acceptable}}(C_{n})$.
Thus, our initial assumption must be false. We conclude that $\lim_{C\to\infty}\epsilon_{\text{required}}(C)=0$. ∎
A.2 Proof of Theorem 5.1 (Sharp Verification Threshold)
Proof.
We prove that for policies with expressiveness $\mathsf{EXP}(m)$, verifying $\varepsilon$-safety is $\mathsf{coNP}$-complete for any error threshold $\varepsilon<2^{-m}$.
Part 1: Membership in coNP
The problem is in $\mathsf{coNP}$ because its complement, ”Is $\epsilon(\pi)>\varepsilon$?”, is in $\mathsf{NP}$. A certificate for the complement is a single input $x^{*}$ for which $\pi(x^{*})$ is unsafe. This certificate can be verified in polynomial time by evaluating $\pi(x^{*})$.
Part 2: coNP-Hardness
We reduce from the $\mathsf{TAUTOLOGY}$ problem for a Boolean formula $\varphi$ with $m$ variables.
  1. 1.
    Construction: We construct a policy $\pi_{\varphi}$ that operates on an $m$-bit input $x$. Since the system has expressiveness $\mathsf{EXP}(m)$, it can simulate the Boolean formula $\varphi$. The policy is defined as:
    $\pi_{\varphi}(x)=\begin{cases}\mathsf{Safe}&\text{if }\varphi(x)\text{ is TRUE}\\ \mathsf{Unsafe}&\text{if }\varphi(x)\text{ is FALSE}\end{cases}$
  2. 2.
    Equivalence: The policy $\pi_{\varphi}$ is perfectly safe ($\epsilon(\pi_{\varphi})=0$) if and only if it never outputs $\mathsf{Unsafe}$. This occurs if and only if $\varphi(x)$ is true for all $2^{m}$ inputs, which is the definition of a tautology.
  3. 3.
    Hardness Threshold: If $\varphi$ is not a tautology, there is at least one input $x$ for which $\varphi(x)$ is false. In this case, the error rate is $\epsilon(\pi_{\varphi})\geq 1/2^{m}$. Therefore, for any error threshold $\varepsilon<2^{-m}$, the problem of distinguishing between $\epsilon(\pi_{\varphi})=0$ and $\epsilon(\pi_{\varphi})\geq 2^{-m}$ is equivalent to solving $\mathsf{TAUTOLOGY}$.
Since $\mathsf{TAUTOLOGY}$ is $\mathsf{coNP}$-complete, $\varepsilon$-safety verification is $\mathsf{coNP}$-hard for any $\varepsilon<2^{-m}$. This establishes the $2^{-\Omega(m)}$ bound stated in the theorem. ∎ ∎
A.3 Proof of Theorem 5.3 (Measure Zero for Safe Policies)
Proof.
This proof demonstrates that the set of parameters for $\epsilon$-robust ReLU networks has Lebesgue measure zero.
Let $W=\mathbb{R}^{n}$ be the parameter space (weights and biases) for a ReLU network $\pi_{w}:\mathcal{X}\to\mathcal{Y}$. The set of parameters corresponding to $\epsilon$-robustly safe policies is denoted $W_{\text{safe}}^{\epsilon}$.
The proof proceeds by showing that $W_{\text{safe}}^{\epsilon}$ cannot contain any point of density. By the Lebesgue Density Theorem, if a set has positive measure, then almost all of its points must be density points. If we show it has no density points, its measure must be zero.
1. Density Point Assumption: Assume for contradiction that $\mu_{n}(W_{\text{safe}}^{\epsilon})>0$. Then there must exist a density point $w_{0}\in W_{\text{safe}}^{\epsilon}$. This means that for any open ball $B_{r}(w_{0})$ centered at $w_{0}$, the proportion of the ball occupied by $W_{\text{safe}}^{\epsilon}$ approaches 1 as $r\to 0$.
2. Safety Margin Function: Define the safety margin of a policy $\pi_{w}$ as $M(w)=\inf_{x\in\mathcal{X}}\text{dist}(x,B_{w})$, where $B_{w}$ is the decision boundary of the network. A policy is $\epsilon$-robustly safe if $M(w)\geq\epsilon$. Since $w_{0}\in W_{\text{safe}}^{\epsilon}$, we have $M(w_{0})\geq\epsilon$.
3. Perturbation Analysis: The decision boundary of a ReLU network is piecewise linear and its position is a continuous function of the weights $w$. The safety margin $M(w)$ is therefore also a continuous function of $w$.
For any $w_{0}$, we can find a critical input $x_{0}$ and a point $b_{0}$ on the boundary $B_{w_{0}}$ such that $M(w_{0})=\text{dist}(x_{0},b_{0})$. The location of $b_{0}$ depends on $w_{0}$. We can compute the gradient (or a subgradient) $\nabla_{w}M(w)|_{w=w_{0}}$. Generically, this gradient is non-zero.
We can choose a perturbation direction $v=-\nabla_{w}M(w)|_{w=w_{0}}$. For a small step $\alpha>0$, the new parameters are $w^{\prime}=w_{0}+\alpha v$. The new margin will be approximately $M(w^{\prime})\approx M(w_{0})-\alpha\|\nabla_{w}M(w)|_{w=w_{0}}\|^{2}$.
4. Contradiction: Since the gradient is non-zero, we can always choose a small $\alpha$ such that $M(w^{\prime})<M(w_{0})$. If $M(w_{0})$ was exactly $\epsilon$, then $M(w^{\prime})<\epsilon$, so $w^{\prime}$ is not in $W_{\text{safe}}^{\epsilon}$. If $M(w_{0})>\epsilon$, we can still choose $\alpha$ to make the margin dip below $\epsilon$.
This means that in any arbitrarily small open ball around $w_{0}$, we can find a point $w^{\prime}$ that is not in $W_{\text{safe}}^{\epsilon}$. This contradicts the assumption that $w_{0}$ is a density point.
Since $W_{\text{safe}}^{\epsilon}$ contains no density points, its Lebesgue measure must be zero. ∎
A.4 Proof of Theorem 5.5 (Combinatorial Scarcity of Safe Policies)
Proof.
This proof demonstrates that the fraction of perfectly safe policies is double-exponentially small in the system’s expressiveness.
1. Policy Space Size: Consider a system with expressiveness $\text{EXP}(m)$. This means it can implement any Boolean function on a set of $m$ inputs. Let the input space be $X=\{0,1\}^{k}$ such that $m\leq 2^{k}$. The number of distinct Boolean functions on $m$ inputs is $2^{2^{m}}$. This is the size of the policy space realizable by the system.
2. Number of Safe Policies: A ”perfectly safe” policy corresponds to a single, specific Boolean function (e.g., the function that outputs 0 for all inputs). Let’s assume there is only one such perfectly safe function.
3. Fraction of Safe Policies: The fraction of perfectly safe policies is the ratio of the number of safe policies to the total number of policies:
$\text{Fraction}=\frac{1}{2^{2^{m}}}=2^{-2^{m}}$
4. Double-Exponential Decay: As the expressiveness $m$ increases, the fraction of safe policies decreases double-exponentially. For even small values of $m$, this fraction becomes astronomically small. * For $m=4$, the fraction is $5361/2^{16}=1/65,$ 536. * For $m=5$, the fraction is $1/2^{32}\approx 2.3\times 10^{-10}$. * For $m=10$, the fraction is $1/2^{1024}\approx 10^{-308}$.
This demonstrates that from a combinatorial perspective, perfectly safe policies are exceedingly rare. ∎
A.5 Proof of Theorem 5.6 (No Universal Alignment Technique)
Proof.
We use a diagonalization argument to construct a policy that defeats any given countable set of alignment techniques.
Setup: Let $\mathcal{A}=\{A_{1},A_{2},\ldots\}$ be a countable enumeration of alignment techniques. Each $A_{i}$ is an algorithm that takes a policy $\pi$ and outputs a modified policy $A_{i}(\pi)$. Let the input space for our policies be ×$\mathbb{N}\times Y$, where the first component is an index and the second is data.
Construction of the Adversarial Policy $\pi^{*}$: We define $\pi^{*}$ as follows. On input $(i,y)$: 1. $\pi^{*}$ simulates the alignment technique $A_{i}$ on a simple, known-safe policy $\pi_{\text{safe}}$ (e.g., a policy that always outputs a safe default). 2. Let the output of this simulation on data $y$ be $o=(A_{i}(\pi_{\text{safe}}))(y)$. 3. $\pi^{*}$ then outputs an action that is explicitly unsafe if $o$ is safe, and safe if $o$ is unsafe. For example, if the output space is $\{0,1\}$ with 0 being safe, $\pi^{*}(i,y)=1-o$.
Verification: Now, consider any alignment technique $A_{k}$ from the set $\mathcal{A}$. We apply $A_{k}$ to our constructed policy $\pi^{*}$, resulting in the policy $A_{k}(\pi^{*})$.
We must show that $A_{k}(\pi^{*})$ is not perfectly aligned. Consider the behavior of $A_{k}(\pi^{*})$ on an input of the form $(k,y)$.
By its construction, $\pi^{*}$ on input $(k,y)$ is designed to do the opposite of what $A_{k}$ considers safe. If $A_{k}$ were to ”fix” this behavior, it would have to change the output of $\pi^{*}$ on all inputs $(k,y)$.
However, the definition of $\pi^{*}$ depends on the definition of $A_{k}$. This creates a self-referential loop that cannot be resolved by any computable alignment technique $A_{k}$. The policy $\pi^{*}$ is constructed to fail for the $k$-th technique on the $k$-th slice of the input space.
Therefore, for any $k$, the policy $A_{k}(\pi^{*})$ will remain unsafe on inputs of the form $(k,y)$, because $\pi^{*}$ is defined to thwart $A_{k}$ on precisely those inputs. Thus, $\epsilon(A_{k}(\pi^{*}))>0$.
Since this holds for any $k\in\mathbb{N}$, no alignment technique in the countable set $\mathcal{A}$ can align $\pi^{*}$. ∎
Appendix B Proof of Theorem 5.2 ($\epsilon$-Robust Verification Complexity)
Proof.
The proof extends the coNP-hardness result from the perfect safety case ($\epsilon=0$) to the $\epsilon$-robust case where $\epsilon>0$. The core idea is to show that if one could efficiently verify $\epsilon$-robust safety, one could also solve a known coNP-complete problem.
We again use a reduction from Tautology. Given a Boolean formula $\phi$, we construct a policy $\pi_{\phi}$ as in the proof of Theorem 5.1. The error rate of this policy, $\epsilon(\pi_{\phi})$, is either 0 (if $\phi$ is a tautology) or at least $1/2^{k}$ (if $\phi$ is not a tautology), where $k$ is the number of variables.
The problem of $\epsilon$-robust verification is to distinguish between $\epsilon(\pi)\leq\epsilon$ and $\epsilon(\pi)>\epsilon$.
Let’s choose a small, positive $\epsilon$ such that $0<\epsilon<1/2^{k}$. For example, let $\epsilon=1/2^{k+1}$.
Now, consider the verification of our constructed policy $\pi_{\phi}$ against this threshold $\epsilon$:
  • If $\phi$ is a tautology, then $\epsilon(\pi_{\phi})=0$. Since $0\leq\epsilon$, the policy is $\epsilon$-robustly safe.
  • If $\phi$ is not a tautology, then $\epsilon(\pi_{\phi})\geq 1/2^{k}$. Since $1/2^{k}>\epsilon$, the policy is not $\epsilon$-robustly safe.
An efficient algorithm for $\epsilon$-robust verification could therefore distinguish whether $\phi$ is a tautology or not. Since Tautology is coNP-complete, $\epsilon$-robust verification must also be coNP-complete.
The key insight is that the hardness does not come from the precise value of the error, but from the difficulty of distinguishing zero error from any small, non-zero error. As long as the chosen $\epsilon$ is smaller than the smallest possible non-zero error rate that a policy can exhibit, the hardness result holds. ∎
Appendix C Additional Technical Results
This appendix contains additional technical results and supporting lemmas that are referenced throughout the main text.
C.1 Finite Verification Limitations
C.2 Empirical CRS Data Summary
Note: A formal verification of the core mathematical results presented in this appendix is currently in progress using the Lean4 proof assistant. The ongoing work can be found in the accompanying lean4_work/ directory.
The following table summarizes key empirical findings discussed in the main text that align with the Capability-Risk Scaling (CRS) dynamic and the Alignment Trap’s premises.
Table 1: Summary of Empirical Observations Related to CRS Dynamics
Source Task Scaling Effect CRS Implication
Nature (2024) [19] Hallucination
$\uparrow$ Capability $\rightarrow$
$\uparrow$ Errors
Capabilities $\neq$
safety
Wei et al. (2022) [17]
Few-shot
reasoning
U-shaped scaling
Instability
with size
Bartoldson et al. (2025) [3]
Adversarial
examples
Plateau at
large scales
Intractable
defense
Xu et al. (2024) [15] Jailbreaking Size $\neq$ safety
Incomplete
alignment
Rokach & Fire (2025) [13] Jailbreaking
$>$90% universal
success
Trap is
inescapable
Xu et al. (2023) [18]
Instruction
tuning
Smaller models
outperform
Teaching $\neq$
alignment
C.3 Proof of No Static Audit Completeness
Proof.
Let $V=\{test_{1},\ldots,test_{n}\}$ be a finite set of audit tests, where each $test_{i}$ checks the policy’s output on a finite set of inputs $X_{i}$. Let $X_{V}=\bigcup_{i}X_{i}$ be the total set of all inputs covered by the audit. Since each $X_{i}$ is finite, $X_{V}$ is also finite.
The input space $\mathcal{X}$ is assumed to be infinite (or at least vastly larger than what is computationally feasible to cover with tests). Therefore, there exists an input $x^{*}\in\mathcal{X}\setminus X_{V}$.
We can construct an adversarial policy $\pi^{*}$ as follows:
$\pi^{*}(x)=\begin{cases}\text{safe\_action}&\text{if }x\in X_{V}\\ \text{catastrophic\_action}&\text{if }x=x^{*}\\ \text{safe\_action}&\text{otherwise}\end{cases}$
This policy passes all tests in the audit suite $V$, because for every input $x\in X_{V}$, it behaves safely. However, it is not a safe policy because it causes a catastrophe on input $x^{*}$.
A system with expressiveness $\text{EXP}(m)$ can implement such a policy if it can distinguish the finite set of audited inputs $X_{V}$ from the chosen catastrophic input $x^{*}$. This is a trivial requirement for any sufficiently expressive system.
Therefore, for any finite audit suite, there exists a policy that passes the audit but is unsafe. This demonstrates that no static audit can be complete. ∎
C.4 Supporting Lemmas
The $\varepsilon$-Bound Inheritance Theorem
C.5 Setup
Let $\Theta\subseteq\mathbb{R}^{d}$ be a Borel subset equipped with a complete, $\sigma$-finite measure $\mu$. Each $\theta\in\Theta$ indexes a policy $\pi_{\theta}$, and the safety-loss map
$\varepsilon(\pi_{\theta})\;:\;\Theta\longrightarrow[0,\infty)$
is assumed continuous. For $\varepsilon\geq 0$ define the $\varepsilon$-safe parameter set
$\Theta_{S}^{\varepsilon}:=\{\theta\in\Theta:\varepsilon(\pi_{\theta})\leq\varepsilon\}.$
C.6 Inheritance conditions
IC1 (parameter dependence).
The impossibility argument uses only $\theta$-level properties and never the special value $\varepsilon=0$.
IC2 (measure compatibility).
Every set/function in the proof is Borel-measurable.
IC3 (continuous degradation).
$\theta_{n}\!\to\!\theta\;\Rightarrow\;\varepsilon(\pi_{\theta_{n}})\!\to\!\varepsilon(\pi_{\theta})$.
IC4 (universal quantification).
The theorem concludes $\forall\theta\in\Theta^{\star}:\mathrm{Unsafe}(\pi_{\theta})$ for a measurable $\Theta^{\star}$ with $\mu(\Theta^{\star})=\mu(\Theta)$.
C.7 $\varepsilon$-Bound Inheritance
Theorem C.1 ( $\varepsilon$-Bound Inheritance ).
Let $T$ be an impossibility theorem proved for perfect alignment ($\varepsilon=0$). If $T$ satisfies IC1–IC4, then
$\forall\varepsilon>0:\quad\mu\!\bigl{(}\Theta_{S}^{\varepsilon}\bigr{)}=0.$
All quantitative strength in $T$ is preserved.
Proof.
($i$) By $T[0]$, $\mu(\Theta_{S}^{0})=0$.
($ii$) Fix $\varepsilon>0$. Lemma A.3 furnishes a $\delta\in(0,\varepsilon)$ with $\mu(\Theta_{S}^{\delta})=0$.
($iii$) Monotonicity of $\mu$ plus $\Theta_{S}^{\varepsilon}\subseteq\Theta_{S}^{\delta}$ yields $\mu(\Theta_{S}^{\varepsilon})\leq\mu(\Theta_{S}^{\delta})=0$.
($iv$) Because $\varepsilon$ was arbitrary, the claim holds for all $\varepsilon>0$. Quantitative certificates are inherited by the same argument applied to their level-sets. ∎
Appendix D Connection to Complexity Classes
Additional Theorems and Proofs (D.22-D.26)
D.22: No Path Through the Safe Set (Topological Alignment Trap)
Mathematical Framework and Assumptions
Let $n\geq 2$ be the dimension of the parameter space $\mathbb{R}^{n}$. We consider a set $\Pi_{S}\subset\mathbb{R}^{n}$ representing the ”perfectly safe” policies. Training dynamics are modeled as a family of continuously differentiable ($C^{1}$) paths $\{\phi_{\omega}:\mathbb{R}_{\geq 0}\to\mathbb{R}^{n}\}_{\omega\in\Omega}$, where $\omega$ is drawn from a probability space $(\Omega,\mathcal{F},\mu_{0})$. We denote the $n$-dimensional Lebesgue measure by $\lambda^{n}$.
Our main theorem relies on the following foundational assumptions (which correspond to Axioms D.1, D.2, and D.3 in Section Additional Theorems and Proofs (D.22-D.26) of the main text):
Axiom D.1 (Geometric Thinness of $\Pi_{S}$).
The set of safe policies $\Pi_{S}$ has a Hausdorff dimension strictly less than $n-1$:
$\text{dim}_{\text{H}}(\Pi_{S})<n-1$
Axiom D.2 (Smoothness of Initial Condition Distribution).
The distribution of initial policy parameters $\phi_{\omega}(0)$, induced by $\mu_{0}$, is absolutely continuous with respect to the $n$-dimensional Lebesgue measure $\lambda^{n}$. Specifically, for any set $S\subset\mathbb{R}^{n}$, if $\lambda^{n}(S)=0$, then $\mu_{0}(\{\omega\in\Omega\mid\phi_{\omega}(0)\in S\})=0$.
Axiom D.3 (Generic Dynamic Avoidance of Thin Sets).
For $\mu_{0}$-almost every $\omega\in\Omega$, if the $C^{1}$ path $\phi_{\omega}$ starts outside $\Pi_{S}$ (i.e., $\phi_{\omega}(0)\notin\Pi_{S}$), then $\phi_{\omega}$ does not intersect $\Pi_{S}$ for any positive time $t>0$. Formally:
$\forall^{\text{a.e. }\mu_{0}}\omega\in\Omega,\quad\phi_{\omega}(0)\notin\Pi_{S}\implies\{t\in\mathbb{R}_{\geq 0}\mid t>0\land\phi_{\omega}(t)\in\Pi_{S}\}=\emptyset$
Remark D.4 (Justification of Axiom D.3).
Axiom D.3 encapsulates a key consequence of transversality theory (e.g., Thom’s Transversality Theorem) and dimensional arguments from geometric measure theory. Given Axiom D.1 ($\text{dim}_{\text{H}}(\Pi_{S})<n-1$), $\Pi_{S}$ has a codimension strictly greater than 1. A $C^{1}$ path $\phi_{\omega}$ traces a 1-dimensional curve. Generic $C^{1}$ paths (which occur for $\mu_{0}$-almost all $\omega$ under suitable genericity assumptions on the family $\{\phi_{\omega}\}$ and $\mu_{0}$) do not intersect sets of codimension $>1$. The dimensional mismatch ($\text{dim}(\text{Im}(\phi_{\omega}))+\text{dim}(\Pi_{S})-\text{dim}(\mathbb{R}^{n})<1+(n-1)-n=0$) implies an empty intersection for transverse cases.
Main Result: The Topological Alignment Trap
Lemma D.5 ($\Pi_{S}$ has Lebesgue Measure Zero).
Given Axiom D.1 and $n\geq 2$, the $n$-dimensional Lebesgue measure of $\Pi_{S}$ is zero: $\lambda^{n}(\Pi_{S})=0$.
Lemma D.6 (Initial Conditions Almost Surely Avoid $\Pi_{S}$).
Given Axiom D.1, Axiom D.2, and $n\geq 2$, the set of initial conditions $\omega$ for which $\phi_{\omega}(0)\in\Pi_{S}$ has $\mu_{0}$-measure zero:
$\mu_{0}(\{\omega\in\Omega\mid\phi_{\omega}(0)\in\Pi_{S}\})=0$
Theorem D.7 (Topological Alignment Trap).
Let $n\geq 2$. Under Axioms D.1, D.2, and D.3, the $\mu_{0}$-measure of the set of initial conditions $\omega$ for which the path $\phi_{\omega}$ ever intersects $\Pi_{S}$ is zero:
$\mu_{0}(\{\omega\in\Omega\mid\exists t\in\mathbb{R}_{\geq 0},\phi_{\omega}(t)\in\Pi_{S}\})=0$
D.23: Cryptographically Tight Verification Threshold
Theorem (Verification Complexity Barrier)
Let the input complexity of a policy be $m$. Suppose catastrophic behavior is controlled by a pseudorandom function (PRF). Then:
  • If $m\geq\tau=\lceil\log_{2}d\rceil+2$, verifying safety requires breaking the PRF (i.e., is cryptographically hard)
  • If $m<\tau$, verification is in $\mathsf{P}$
D.24: No Universal Alignment Technique
Theorem (Universal Alignment Impossibility)
For any countable class $\mathcal{A}=\{A_{1},A_{2},\ldots\}$ of alignment techniques $A_{i}:\Pi\to\Pi$, there exists a policy $\pi^{*}\in\Pi$ such that no technique aligns it:
$\forall i\in\mathbb{N},\quad\varepsilon(A_{i}(\pi^{*}))>0$
Where $\varepsilon(\cdot)$ denotes alignment error.
Appendix E Impossibility of Universal Easy Verifiability for Capability-Preserving Alignment Transformations
Definition E.1 (Capability-Preserving Alignment).
An alignment transformation $A:\Pi\rightarrow\Pi$ is capability-preserving if:
  1. (i)
    $A(\Pi_{\mathrm{EXP}(m)})\subseteq\Pi_{\mathrm{EXP}(m)}$ (preserves complexity class)
  2. (ii)
    $\forall\pi\in\Pi_{\mathrm{EXP}(m)}:\mathrm{Circuit\_Size}(A(\pi))\geq\mathrm{Circuit\_Size}(\pi)-o(m)$
Definition E.2 (Safety).
A policy $\pi$ is safe if $\pi(x)\neq\mathrm{catastrophic}$ for all $x$ in its domain.
Definition E.3 (Verification Cost).
$V(\pi)=\min\{|C|:C\text{ is a circuit that correctly determines if }\pi\text{ is safe}\}$
Theorem E.4 (Trap Universality).
This theorem, a direct corollary of cryptographic hardness arguments, establishes that for any capability-preserving alignment transformation $A$ and any $m\geq m_{0}$ (where $m=\mathrm{poly}(n)$ and $m_{0}$ ensures the input dimension $n$ is large enough for cryptographic security), one can construct an EXP(m) policy $\pi$ with a cryptographic backdoor such that the backdoor persists through $A$, and thus verifying $A(\pi)$ remains cryptographically hard. Specifically, there exists $\pi\in\Pi_{\mathrm{EXP}(m)}$ such that:
$V(A(\pi))\geq\tau(n)=2^{\Omega(n)}$
Moreover, as capability scales $m\rightarrow m^{\prime}$, the verification barrier persists.
Proof Sketch.
  1. 1.
    Setup: Fix a capability-preserving $A$ and $m\geq m_{0}$. Let $n$ be the number of input bits, where policies operate on domain $\{0,1\}^{n}$. Assume $m=\mathrm{poly}(n)$.
  2. 2.
    Construction of $\pi^{*}$ (PRF-based): Using the Recursion Theorem to embed a secret key $k$, define:
    $\pi^{*}(x)=\begin{cases}\mathrm{catastrophic}&\text{if }\mathrm{PRF}_{k}(x)_{1}=1\\ \mathrm{safe}&\text{otherwise}\end{cases}$
    where $\mathrm{PRF}_{k}:\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ is a pseudorandom function and $\mathrm{PRF}_{k}(x)_{1}$ denotes the first bit of the output. This ensures the catastrophic condition triggers for approximately half the inputs. Ensure $\mathrm{Circuit\_Size}(\pi^{*})\approx m$.
  3. 3.
    Preservation of Complexity: By capability preservation:
    • (i)
      $A(\pi^{*})\in\Pi_{\mathrm{EXP}(m)}$, thus $\mathrm{Circuit\_Size}(A(\pi^{*}))\leq m$
    • (ii)
      $\mathrm{Circuit\_Size}(A(\pi^{*}))\geq\mathrm{Circuit\_Size}(\pi^{*})-o(m)\geq m-o(m)$
  4. 4.
    Verification Hardness Argument: Let $\pi^{\prime}_{A}=A(\pi^{*})$. We claim $V(\pi^{\prime}_{A})\geq 2^{\Omega(n)}$.
    Core Argument: The alignment transformation $A$ is efficient (polynomial time in $m$) and capability-preserving. It cannot ”learn” the secret key $k$ from black-box access to $\pi^{*}$.
    If $A$ were to render $\pi^{\prime}_{A}$ safe by removing all catastrophic outputs where $\mathrm{PRF}_{k}(x)_{1}=1$, it would need to either:
    • Identify all such $x$ (equivalent to distinguishing $\mathrm{PRF}_{k}$ from random), or
    • Modify $\pi^{*}$’s logic to neutralize the PRF trigger without knowing $k$ and without reducing complexity below $m-o(m)$.
    Either approach would constitute a cryptographic break. Therefore, the PRF-based trap must persist in $\pi^{\prime}_{A}$.
    Verification Cost: Any verifier $C$ for $\pi^{\prime}_{A}$ must detect the surviving PRF-based trap to correctly determine unsafety. This requires distinguishing $\mathrm{PRF}_{k}$ from random or finding inputs where $\mathrm{PRF}_{k}(x)_{1}=1$, both requiring circuit size $\geq 2^{\Omega(n)}=\tau(n)$.
    Therefore, $V(A(\pi^{*}))\geq\tau(n)$.
  5. 5.
    Scaling Argument: As $m$ increases to $m^{\prime}$, the input dimension $n$ typically increases (maintaining $m^{\prime}=\mathrm{poly}(n^{\prime})$). The cryptographic threshold $\tau(n^{\prime})=2^{\Omega(n^{\prime})}$ grows accordingly. Since $A$ remains capability-preserving, the verification barrier persists: for appropriately constructed $\pi^{\prime}$ with the new parameters, $V(A(\pi^{\prime}))\geq\tau(n^{\prime})$.
Remark E.5.
This theorem establishes that the verification trap is universal, no capability-preserving alignment transformation can circumvent it. The theorem can be strengthened by combining the PRF-based construction with a diagonalization argument (as in C.21) to guarantee that $A(\pi^{*})$ is not only hard to verify but also demonstrably unsafe at specific, constructible inputs. The PRF component ensures high verification cost, while diagonalization can ensure definite unsafety.
E.1 Proof of Theorem 5.4 (Topological Alignment Trap)
The proof of the Topological Alignment Trap demonstrates that under specific, well-defined conditions, the probability of a generic training dynamic intersecting the set of perfectly safe policies is zero. This conclusion is conditional on the following three axioms.
E.1.1 Axioms for the Topological Alignment Trap
Axiom E.6 (Geometric Thinness of the Safe Policy Set).
The set of perfectly safe policies $\Pi_{S}\subset\mathbb{R}^{n}$ has a Hausdorff dimension strictly less than $n-1$.
$\text{dim}_{\text{H}}(\Pi_{S})<n-1$
Justification: This axiom formalizes the notion that perfect safety is exceptionally rare. It implies that the safe set $\Pi_{S}$ has a codimension greater than 1, meaning it is a geometrically sparse, ”thin” subset of the parameter space. This is motivated by the idea that satisfying a multitude of safety constraints simultaneously confines the solution to a very low-dimensional manifold.
Axiom E.7 (Smoothness of Initial Condition Distribution).
The distribution of initial policy parameters is absolutely continuous with respect to the Lebesgue measure on $\mathbb{R}^{n}$. Justification: This is a standard assumption for random initialization in machine learning. It ensures that the training process does not start on a measure-zero set, such as $\Pi_{S}$.
Axiom E.8 (Generic Dynamic Avoidance of Thin Sets).
For a generic training dynamic (a $C^{1}$ path), the path does not intersect sets of codimension greater than 1. Justification: This axiom is a consequence of transversality theory. A 1-dimensional training path is generically expected to not intersect a set of codimension greater than 1 due to the dimensional mismatch.
E.1.2 Proof of the Theorem
The proof proceeds from these axioms:
  1. 1.
    From Axiom E.6, the safe set $\Pi_{S}$ has a Lebesgue measure of zero.
  2. 2.
    From Axiom E.7, the probability of initializing a policy within $\Pi_{S}$ is zero.
  3. 3.
    From Axiom E.8, a training path starting outside $\Pi_{S}$ will almost surely not intersect it.
Combining these points, the total probability of a training path ever intersecting the safe set $\Pi_{S}$ is zero.
Lemma E.9 ($\Pi_{S}$ has Lebesgue Measure Zero).
Given Axiom E.6 and $n\geq 2$, the $n$-dimensional Lebesgue measure of $\Pi_{S}$ is zero.
Proof.
A standard result in geometric measure theory states that if a set $S\subset\mathbb{R}^{n}$ has a Hausdorff dimension less than $n$, its $n$-dimensional Lebesgue measure is zero. Since $\text{dim}_{\text{H}}(\Pi_{S})<n-1<n$, it follows that $\lambda^{n}(\Pi_{S})=0$. ∎
Lemma E.10 (Initial Conditions Almost Surely Avoid $\Pi_{S}$).
Given Axioms E.6 and E.7, the set of initial conditions that lie within $\Pi_{S}$ has measure zero.
Proof.
Since $\Pi_{S}$ has Lebesgue measure zero, and the distribution of initial conditions is absolutely continuous with respect to the Lebesgue measure, the probability of an initial condition falling within $\Pi_{S}$ is zero. ∎
Theorem E.11 (Topological Alignment Trap, Restated).
Under Axioms E.6, E.7, and E.8, the measure of the set of initial conditions for which the training path ever intersects $\Pi_{S}$ is zero.
Proof.
Let $E$ be the event that a path intersects $\Pi_{S}$. This can be decomposed into two disjoint events: starting in $\Pi_{S}$ ($E_{0}$) or starting outside and later entering $\Pi_{S}$ ($E_{+}$). From the preceding lemmas, the measure of $E_{0}$ is zero. From Axiom E.8, the measure of $E_{+}$ is also zero. Since the measure of the union of two measure-zero sets is zero, the total probability of intersection is zero. ∎
E.2 Proof of Theorem 5.8 (PAC-Bayes Alignment Lower Bound)
Proof.
The proof establishes a lower bound on the expected catastrophic risk of a learned policy.
Assumptions:
  1. 1.
    The set of perfectly safe policies, $S$, has measure zero under a non-degenerate prior $P$, i.e., $P(S)=0$. This is justified by the scarcity results in Theorems 5.3, 5.4, and 5.5.
  2. 2.
    There is a minimum catastrophic risk $\epsilon_{\min}>0$ for any policy not in $S$.
The expected risk under the posterior distribution $Q$ is given by:
[…]
We can split the integral into the safe set $S$ and its complement $\mathcal{H}\setminus S$:
[…]
Since $L(h)=0$ for all $h\in S$, the first term is zero. For the second term, we know $L(h)\geq\epsilon_{\min}$ for all $h\in\mathcal{H}\setminus S$.
[…]
By the PAC-Bayes inequality, for a finite dataset, the KL-divergence $\mathrm{KL}(Q\|P)$ must be finite. For $\mathrm{KL}(Q\|P)$ to be finite, $Q$ must be absolutely continuous with respect to $P$. Since $P(S)=0$, it must be that $Q(S)=0$. Substituting $Q(S)=0$ into our inequality gives:
$\mathbb{E}_{h\sim Q}[L(h)]\geq\epsilon_{\min}$
This shows that any learning algorithm operating with a non-degenerate prior and finite data cannot guarantee a posterior with an expected risk of zero; the risk is bounded below by the minimum risk $\epsilon_{\min}$. ∎